OpenStudy (anonymous):

Danny has decided that he's sick and tired of the cucumbers and tomatoes he buys at the local supermarket. This year, he's going to grow some of his own. However, he is worried about rabbits because they like to eat young plants. Fortunately, he has 60 feet of wire fencing to put around his garden. Since he wants to plant nice neat rows, he wants his garden to be rectangular. He's not sure exactly what size he should make his garden, so he turns to his old friend, algebra. Remember that the perimeter of a rectangle can be found with the formula P=2l+2w . Write an inequality that will give the possible lengths and widths of his rectangular garden so that the perimeter of the garden will not be more than 60. Use x for length of the garden and y for the width.

OpenStudy (ashleyisakitty):

my goodness your school puts effort into their word problem

OpenStudy (anonymous):

I know u hate math but if u draw a picture from the question can answer your self :D

OpenStudy (anonymous):

PLEASE help... :( I have no clue what to do!

OpenStudy (anonymous):

ammm ... your question is tooooo long can u write it shorter ????

OpenStudy (anonymous):

ok... Well write an inequality that will give the possible lengths and widths of his rectangular garden so that the perimeter of the garden will not be more than 60. Use x for the length and y for the width of the garden.

OpenStudy (anonymous):

oK ! do u want learn it too Modom ?:D

OpenStudy (anonymous):

huh ??? yes or no ????

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Very good :) I thought u hate math :)

OpenStudy (anonymous):

hah I DO

OpenStudy (anonymous):

No problem .... Now start with one question : what is the formula to find the perimeter of rectangul? Do u know that ????:D

OpenStudy (anonymous):

Yes, I sure do. P=2l=2w

OpenStudy (anonymous):

OK ! We can write : 2(w+l)too ? OK?

OpenStudy (anonymous):

oh woops. lol its P=2l+2w

OpenStudy (anonymous):

and yes you can

OpenStudy (anonymous):

Thank u :) Now wait ... amm ... Your question want an inequality for perimeter and 60 ??

OpenStudy (anonymous):

it asks for an inequality that will give the possible lengths and widths of his rectangular garden so that the perimeter will not be more than 60. (using x as length and y for width)

OpenStudy (anonymous):

OK ... If we get w=x L=y we have :2(x+Y) right ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So we want it was <60 OK?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

So we had :2(x+y) and <60 So it can = 2(x+y) too OK ? So we have : 2(x+y)=<60 Got it ?

OpenStudy (anonymous):

yes

hero (hero):

"Your question is too long can you write it shorter" LOL @E.ali

hero (hero):

I don't think @IloveCharlie had any choice in the matter. A student couldn't ask the teacher to write the question shorter if it were on a test.

OpenStudy (anonymous):

amm ... Your mean ??? @IloveCharlie :that was answer :) Just it :)

OpenStudy (anonymous):

Wait WHAT?! THATS IT?!

OpenStudy (anonymous):

Yep !

OpenStudy (anonymous):

2(x+y)<60 THATS IT!?

OpenStudy (anonymous):

yes !

OpenStudy (anonymous):

wow

hero (hero):

Actually, you're supposed to express the distinct inequalities for both the length and the width.

OpenStudy (anonymous):

but it can =60 too :)

hero (hero):

something like l < something w < something

OpenStudy (anonymous):

So the < would be underlined?

OpenStudy (anonymous):

@Hero, so im not done? lol theres more to it?

hero (hero):

Yes, I'm sorry to have to be the one to tell you.

OpenStudy (anonymous):

aha ! If it was look : w+L=30 so if w>L we have another ...

OpenStudy (anonymous):

huh im confused lol. So what is the REAL answer?

OpenStudy (anonymous):

Please wait ...

hero (hero):

But all you do is start with \(2x + 2y \le 60\) Then isolate x for length Then isolate y for width

OpenStudy (anonymous):

k lemme try @Hero gimme a second, don't leave. Im a lil slow with math haha

OpenStudy (anonymous):

yeeeah. Im not copying. lol. How do I isolate x and y?

OpenStudy (anonymous):

get them in w&l Just it :)

hero (hero):

Well, I will isolate x step by step. Then afterwards, you can try isolating y.

OpenStudy (anonymous):

how lol :/

OpenStudy (anonymous):

Look :